# Fault Tolerant Zero Forcing

### v2

@@ 1 Introduction
Zero forcing on a graph was introduced in [AIM2008] to bound the maximum nullity obtained in space of matrices defined by the graph. We delay formal definitions until Section 2, but informally the goal of zero forcing is to determine the smallest size of a set of vertices that when these vertices are the only vertices initially colored blue, the remaining vertices all become blue after a finite number of applications of the zero forcing color change rule.
Since its introduction, there has been significant interest in zero forcing and a host of related parameters that are defined using a similar paradigm. Though first studied for its applications to the problem of determining the minimum rank of a symmetric matrix associated with a graph [AIM2008, allison2010minimum, parameters2013, ekstrand2013psdzf], zero forcing and its variants have also been studied for their applications to quantum control [quantum], the spread of disease and information [DREYER2009], and monitoring electric grids[haynes2002domination]. The latter application is related to the placement of phase-monitoring units (PMUs) in an electrical power network, and thus it has motivated investigation into variations of the parameter when PMUs are allowed to fail [bjorkman2023robust, pai2010restricted].
The property of robustness when searching a graph is a desirable property, and in this article we introduce a fault tolerant variant of zero forcing and study its properties. While the leaky-forcing variant of zero forcing studied in [abbas2025, alameda2024leaky, alameda2023gen] seeks to provide robustness of the zero forcing process when certain vertices are not allowed to contribute to the zero forcing process, our parameter is determined by the behavior of all subsets of a certain size within a prescribed set of vertices. Thus our parameter is novel and distinct.
In [KENTER2019124], zero forcing is related to the solving of linear systems, and leaky forcing was defined as a parameter to study the resilience of solving linear or physical systems where a leak at a node prevents the information available at that from being able to be used reliably in deducing state of other vertices in the system. In contrast, the fault tolerant zero forcing parameter studied herein corresponds to a loss of information at a node (or nodes) but not the reliability of information about the node through the rest of the system.
Much is known about the zero forcing number, including the values of the parameter for many families of graphs, as well as its behavior under various graphs operations, see [hogben2022book] for a detailed account. Additionally, other parameters of a graph derived directly from the zero forcing number and the zero forcing color change rule have been investigated. Examples include the propagation time [hogben2012prop] of a graph and the throttling number [carlson2019throttling] of a graph. We thus have many tools available for studying the fault tolerant zero forcing parameter.
@@ 2 Preliminaries
Let $G$ be a simple graph (i.e. undirected with no loops or multiple edges) and denote its set of vertices by $V(G)$ and its edges by $E(G)$. We denote by $uv$ the edge of $G$ whose endpoints are the vertices $u,v\in V(G)$, while $\{u,v\}$ denotes the subset of $V(G)$ with these vertices. For $V\subseteq V(G)$, we denote the subgraph of $G$ induced by $V$ by $G[V]$. We follow the notation in [hogben2022book], and in particular for a graph with $n$ vertices, we denote path graphs by $P_{n}$, cycle graphs by $C_{n}$, and complete graphs by $K_{n}$. We denote the degree of a vertex $v\in V(G)$ by $\deg_{G}(v)$ or $\deg(v)$ if the context is clear. A vertex $v$ is called universal if $\deg(v)=n-1$, and $v$ is called a leaf if $\deg(v)=1$. Wheels graphs $W_{n}$ are constructed by adding a universal vertex to $C_{n-1}$. The symbols $N(v)$ and $N[v]$ denote the open and closed neighborhoods of $v$, respectively.
@@
A zero forcing set can be thought of as a set of vertices such that if that status of the vertices in the set are known, then the status of every vertex in the graph is determined. For instance, in a space of matrices described by the graph, if the values of entries in a null vector corresponding to a zero forcing set are chosen to be zero, the null vector must be the zero vector. See [hogben2022book] for full details.
@@
Let $G$ be a graph. The $k$-fault tolerant zero forcing number of $G$ is the minimal cardinality of a $k$-fault tolerant zero forcing set for $G$. i.e.
@@
If $G$ is a graph, its connected components are $G_{1},G_{2},\dots,G_{r}$, and $\operatorname{Z}_{t}(G_{i})$ exists for $i=1,2,\dots,r$, then
$\operatorname{Z}_{t}(G)=\displaystyle\sum_{i=1}^{r}\operatorname{Z}_{t}(G_{i}).$
@@
Let $G$ be a graph and suppose $\operatorname{Z}_{t}^{k}(G)$ exists. Then $\operatorname{Z}_{t}^{k}(G)\geq\operatorname{Z}(G)+k$.
@@
For every graph $G$, the standard zero forcing number $\operatorname{Z}(G)$ always exists. However, when fault tolerance is enforced, the existence of $\operatorname{Z}_{t}^{k}(G)$ is not guaranteed. Observe that when a graph $G$ has at least one edge, then at least one force is possible and $\operatorname{Z}(G)\leq n-1$. The existence of $\operatorname{Z}_{t}(G)$ is thus easy to characterize.
Observation 2.8.
@@
The existence of $\operatorname{Z}_{t}^{k}(G)$ for $k\geq 2$ is harder to characterize. Clearly if $G$ is a graph or order $n$ and $\operatorname{Z}(G)+k>n$, then $\operatorname{Z}_{t}^{k}(G)$ does not exist. This tells us immediately that $\operatorname{Z}_{t}^{k}(K_{n})$ does not exist for $k\geq 2$ and any $n\geq 1$, since for $n\geq 2$ we have $\operatorname{Z}(K_{n})=n-1$.
The $k$-fault tolerant zero forcing number can fail to exist for other reasons, however. Let $G$ be any graph, choose a vertex $v\in V(G)$ and attach $2$ pendant vertices, $u$ and $w$, to $v$. If we denote the graph obtained in this manner by $H$, then $\operatorname{Z}_{t}^{k}(H)$ does not exist for $k\geq 2$, since $v$ is the only vertex that can force $u$ or $w$. A more general result is given in Theorem 5.12.
@@ 3 Graph Families
We now establish values of the fault tolerant zero forcing number for graphs which belong to a common family of graphs. The values of the zero forcing number for these graph families were established in [AIM2008].
@@
Let $G=P_{n}$ be a path on $n\geq 2$ vertices. Then $\operatorname{Z}_{t}(G)=2$.
@@
Let $G$ be a path on $n\geq 2$ vertices. Then $\operatorname{Z}(G)=1$ and by the 2.5 we have $\operatorname{Z}_{t}(G)\geq\operatorname{Z}(G)+1=2$. By choosing $B$ to consist of both endpoints, we are guaranteed that every subset of size 1 of $B$ is a zero forcing set, and thus $B$ is a fault tolerant zero forcing set. Since this achieves the lower bound, we have $\operatorname{Z}_{t}(G)=2$. ∎
@@
Let $G=K_{n}$ for $n\geq 2$. Then $\operatorname{Z}_{t}(G)=n$, and $\operatorname{Z}_{t}^{k}(G)$ does not exist for $k\geq 2$.
@@
Let $G=K_{n}$ for some $n\geq 2$. Then $\operatorname{Z}(K_{n})=n-1$ for all $n\geq 2$, and any set of n1n-1 vertices is a minimum zero forcing set. Therefore, the set of all $n$ vertices is a fault tolerant zero forcing set, and thus $\operatorname{Z}_{t}(K_{n})\leq n$. We have $\operatorname{Z}_{t}(K_{n})\geq n$ by 2.5. That $\operatorname{Z}_{t}^{k}(G)$ does not exist for $k\geq 2$ is immediate from 2.5. ∎
@@
$\operatorname{Z}(K_{1,n})=n-1$, and all minimum zero forcing sets are composed of all but one leaf. Thus, the set of all leaves is a 1-fault tolerant set, so $\operatorname{Z}_{t}(K_{1,n})\leq n$. We have $\operatorname{Z}_{t}(K_{1,n})\geq n$ by 2.5. ∎
@@
Let $G=W_{n}$ be a wheel on $n\geq 6$ vertices. Then $\operatorname{Z}_{t}(G)=5$.
@@
Let $G=W_{n}$ for $n\geq 6$. It was shown in [Fallat2014MinimumRM] that $\operatorname{Z}(G)=3$. Furthermore, up to isomorphism there are two minimum zero forcing sets, and these are precisely the minimal zero forcing sets. The first minimum zero forcing set is $B_{1}=\{v_{1},v_{2},v_{3}\}$, where $\deg(v_{1})=\deg(v_{2})=3$, $v_{1}v_{2}\in E(G)$, and $\deg(v_{3})=n-1$, while the second is $B_{2}=\{w_{1},w_{2},w_{3}\}$, where $\deg(w_{1})=\deg(w_{2})=\deg(w_{3})=3$ and $w_{1}w_{2},w_{2}w_{3}\in E(G)$. By 2.5, we have $\operatorname{Z}_{t}(G)\geq 4$. We will show there is no $B$ such that $|B|=4$ is a fault tolerant zero forcing set and construct a fault tolerant zero forcing set $B$ such that $|B|=5$.
@@ 4.1 Compatible path covers
Let $G$ be a graph and $\mathcal{P}=\{P_{i}\}_{i=1}^{r}$ a path cover of $G$. Define $E(\mathcal{P})$ to be the endpoints of paths in $\mathcal{P}$, that is $E(\mathcal{P})=\{v\in V(T)\ |\ \deg_{\mathcal{P}}(v)\leq 1\}$.
@@
Let $G$ be a graph and $v\in V(G)$. A set $\mathfrak{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},\dots,\mathcal{P}_{r}\}$ of path covers of $G$ is called compatible for $S\subseteq V(G)$ if for every $v\in S$:
@@
    either $\deg_{\mathcal{P}_{i}}(v)\geq 2$ for all $i\in\{1,2,\dots,r\}$, or for each $u_{i}\in E(\mathcal{P}_{i})$ such that $\deg_{\mathcal{P}_{i}}(v)=1$, there exists $\mathcal{P}_{j}$ such that $\deg_{\mathcal{P}_{j}}(v)=2$ or $\mathrm{comp}_{\mathcal{P}_{j}}(u_{i})\neq\mathrm{comp}_{\mathcal{P}_{j}}(v)$ and $\deg_{\mathcal{P}_{j}}(u_{i})\geq 1$
We begin with a trivial example to show how compatible collections generalize path covers, before exhibiting a more interesting case.
@@
Consider the Graph $E_{1}$ in Figure 4.1. The collection $\mathfrak{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},\mathcal{P}_{3},\mathcal{P}_{4}\}$, where each $\mathcal{P}_{i}$ is given in Figure 4.2, is a set of path covers compatible for $\{3,4,5,6\}$.
@@
Given a set $\mathfrak{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},\dots,\mathcal{P}_{r}\}$ of path covers of a graph $G$, let $E(\mathfrak{P})=\bigcup_{i=1}^{r}E(\mathcal{P}_{i})$. If $\mathfrak{P}$ is compatible for $S$, then the irredundant vertices of $\mathfrak{P}$ are $I(\mathfrak{P})=E(\mathfrak{P})\setminus S$.
A collections of path covers $\mathfrak{P}$ of a graph $G$ is a minimum compatible collection for $G$ if $|I(\mathfrak{P})|\leq|I(\mathfrak{Q})|$ for every collection of path covers $\mathfrak{Q}$ of $G$. The cardinality of a minimum compatible collection for $G$ is called the irredundance number of $G$ is denoted by $\mathrm{I}(G)$.
@@
Consider $w\not\in B$. Since $w$ must be forced, we have $\deg_{\mathcal{P}_{v}}(w)\geq 1$ for every $v\in B$. If $w\not\in E(\mathfrak{P})$, then $\deg_{\mathcal{P}_{v}}(w)\geq 2$ for all $v\in B$ and $\mathfrak{P}$ is compatible for $w$. So suppose $w\in E(\mathfrak{P})$. For every $\mathcal{P}_{v}$ such that $\deg_{\mathcal{P}_{v}}(w)=1$, the other endpoint of $\mathrm{comp}_{\mathcal{P}_{v}}(w)$ must be in $B\setminus\{v\}$, and thus $\mathfrak{P}$ is compatible for $w$. Therefore $\mathfrak{P}$ is compatible for every $w\not\in B$, and since $I(\mathfrak{P})=|B|$, we have the result. ∎
@@
In general, determining whether a set of compatible path covers for a graph $G$ is minimum is difficult and is a topic that requires further study. All path covers can be enumerated and thus an algorithm exists to test minimality, but this is not effiecient to implement. Some simplifications exist, however, and as an example we offer the next result.
@@
If $\mathfrak{P}$ is a minimum compatible collection for $G$ and $v\in E(\mathfrak{P})$ is isolated in every path cover of $\mathfrak{P}$, then $\mathfrak{P}-v$ is a minimum compatible collection of GvG-v. In particular, under these hypotheses I(G)=I(Gv)+1I(G)=I(G-v)+1

### v3

@@ 1 Introduction
Zero forcing on a graph was introduced in [2] to bound the maximum nullity obtained in space of matrices defined by the graph. We delay formal definitions until Section 2, but informally the goal of zero forcing is to determine the smallest size of a set of vertices that when these vertices are the only vertices initially colored blue, the remaining vertices all become blue after a finite number of applications of the zero forcing color change rule.
Since its introduction, there has been significant interest in zero forcing and a host of related parameters that are defined using a similar paradigm. Though first studied for its applications to the problem of determining the minimum rank of a symmetric matrix associated with a graph [2, 5, 8, 15], zero forcing and its variants have also been studied for their applications to quantum control [11], the spread of disease and information [13], and monitoring electric grids[19]. The latter application is related to the placement of phase-monitoring units (PMUs) in an electrical power network, and thus it has motivated investigation into variations of the parameter when PMUs are allowed to fail [9, 25].
The property of robustness when searching a graph is a desirable property, and in this article we introduce a fault tolerant variant of zero forcing and study its properties. While the leaky-forcing variant of zero forcing studied in [1, 3, 4] seeks to provide robustness of the zero forcing process when certain vertices are not allowed to contribute to the zero forcing process, our parameter is determined by the behavior of all subsets of a certain size within a prescribed set of vertices. Thus our parameter is novel and distinct.
In [23], zero forcing is related to the solving of linear systems, and leaky forcing was defined as a parameter to study the resilience of solving linear or physical systems where a leak at a node prevents the information available at that from being able to be used reliably in deducing state of other vertices in the system. In contrast, the fault tolerant zero forcing parameter studied herein corresponds to a loss of information at a node (or nodes) but not the reliability of information about the node through the rest of the system.
Much is known about the zero forcing number, including the values of the parameter for many families of graphs, as well as its behavior under various graphs operations, see [21] for a detailed account. Additionally, other parameters of a graph derived directly from the zero forcing number and the zero forcing color change rule have been investigated. Examples include the propagation time [20] of a graph and the throttling number [12] of a graph. We thus have many tools available for studying the fault tolerant zero forcing parameter.
@@ 2 Preliminaries
Let $G$ be a simple graph (i.e. undirected with no loops or multiple edges) and denote its set of vertices by $V(G)$ and its edges by $E(G)$. We denote by $uv$ the edge of $G$ whose endpoints are the vertices $u,v\in V(G)$, while $\{u,v\}$ denotes the subset of $V(G)$ with these vertices. For $V\subseteq V(G)$, we denote the subgraph of $G$ induced by $V$ by $G[V]$. We follow the notation in [21], and in particular for a graph with $n$ vertices, we denote path graphs by $P_{n}$, cycle graphs by $C_{n}$, and complete graphs by $K_{n}$. The complete bipartite graph with partite sets of $m$ and $n$ is denoted $K_{m,n}$. We denote the degree of a vertex $v\in V(G)$ by $\deg_{G}(v)$ or $\deg(v)$ if the context is clear. A vertex $v$ is called universal if $\deg(v)=n-1$, and $v$ is called a leaf if $\deg(v)=1$. Wheels graphs $W_{n}$ are constructed by adding a universal vertex to $C_{n-1}$. The symbols $N(v)$ and $N[v]$ denote the open and closed neighborhoods of $v$, respectively. Two vertices $u$ and $v$ of $G$ are twins if $N_{G}(u)=N_{G}(v)$.
@@
A zero forcing set can be thought of as a set of vertices such that if that status of the vertices in the set are known, then the status of every vertex in the graph is determined. For instance, in a space of matrices described by the graph, if the values of entries in a null vector corresponding to a zero forcing set are chosen to be zero, the null vector must be the zero vector. See [21] for full details.
@@
Let $G$ be a graph. The $k$-fault tolerant zero forcing number of $G$ is the minimum cardinality of a $k$-fault tolerant zero forcing set for $G$. i.e.
@@
If $G$ is a graph, its connected components are $G_{1},G_{2},\dots,G_{r}$, and $\operatorname{Z}_{t}^{k}(G_{i})$ exists for $i=1,2,\dots,r$, then
$\operatorname{Z}_{t}^{k}(G)=\displaystyle\sum_{i=1}^{r}\operatorname{Z}_{t}^{k}(G_{i}).$
@@
Let $G$ be a graph and suppose $\operatorname{Z}_{t}^{k}(G)$ exists. Then $\operatorname{Z}_{t}^{k}(G)\geq\operatorname{Z}(G)+k$. Moreover,
$\operatorname{Z}_{t}^{k}(G)>\operatorname{Z}_{t}^{k-1}(G)>\dots>\operatorname{Z}_{t}(G)>\operatorname{Z}(G).$
@@
The study of zero forcing and its variants has been augmented by the introduction of the notion of forts, which are obstructions to zero forcing sets. Specifically, a fort is a nonempty subset $F\subseteq V(G)$ such that for all $v\notin F,|N_{G}(v)\cap F|\neq 1$. Intuitively, a fort is a set of unfilled vertices such that when its complement if filled, no more forces are possible.
Theorem 2.8 ([10],[17]).
Let $G$ be a graph. A set of vertices $S\subseteq V(G)$ is a zero forcing set if and only if $S$ intersects every fort of $G$ nontrivially.
Theorem 2.8 extends naturally to the vertex fault tolerant case.
Observation 2.9.
A set of vertices $S\subseteq V(G)$ is a fault tolerant zero forcing set if and only if $S\cap F\geq k$ for every fort $F$ of $G$.
Proposition 2.10.
Let $G$ be a graph. Then:
  1. 1.
    The only forts of $G$ of cardinality 1 are isolated vertices.
  2. 2.
    The forts of $G$ of cardinality 2 are precisely the sets of (adjacent or independent) twins.
Proof.
Item 1 follows from the fact if $\{v\}$ is a fort, then any neighbor $u$ of $v$ satisfies $|N_{G}(u)\cap\{v\}|=1$, and therefore $v$ has no neighbors. For Item 2, it is clear that a pair of twins is a fort of cardinality 2. If $\{u,v\}$ is fort, then $N_{G}(u)=N_{G}(v)$ since any vertex $w$ in one neighborhood but not the other would satisfy $|N_{G}(w)\cap\{u,v\}|=1$. ∎
For every graph $G$, the standard zero forcing number $\operatorname{Z}(G)$ always exists. However, when fault tolerance is enforced, the existence of $\operatorname{Z}_{t}^{k}(G)$ is not guaranteed. Proposition 2.10 and 2.9 allow us to characterize the existence of $\operatorname{Z}_{t}^{k}(G)$ for $k\leq 2$.
Theorem 2.11.
Let $G$ be a graph. Then:
  1. 1.
@@
  2. 2.
    $\operatorname{Z}_{t}^{2}(G)$ exists if and only if $G$ has neither isolated vertices nor twins.
  3. 3.
    If $u,v$ are twins of $G$, then $\{u,v\}$ is a subset of every fault tolerant zero forcing set.
Note that in particular, Theorem 2.11 implies that if $k\geq 2$, then $\operatorname{Z}_{t}^{k}(G)$ does not exist if $G$ is a complete graph or complete bipartite graph.
@@ 3 Graph Families
We now establish values of the fault tolerant zero forcing number for graphs which belong to a common family of graphs. The values of the zero forcing number for these graph families were established in [2].
@@
Let $G=P_{n}$ be a path on $n\geq 2$ vertices. Then $\operatorname{Z}_{t}(G)=2$. If $n\geq 4$, then $\operatorname{Z}_{t}^{2}(G)=4$.
@@
Let $G$ be a path on $n\geq 2$ vertices. Then $\operatorname{Z}(G)=1$ and by the 2.5 we have $\operatorname{Z}_{t}(G)\geq\operatorname{Z}(G)+1=2$. By choosing $B$ to consist of both endpoints, we are guaranteed that every subset of size 1 of $B$ is a zero forcing set, and thus $B$ is a fault tolerant zero forcing set. Since this achieves the lower bound, we have $\operatorname{Z}_{t}(G)=2$.
Now suppose $n\geq 4$. The graph $G$ contains no 2-fault tolerant sets of size 3, since if the two vertices of degree 1 are white there are no zero forcing sets of size 1. Clearly if $B$ consists of a pair of adjacent degree 2 vertices and the degree 1 vertices, then $B$ is a 2-fault tolerant set since it contains 3 disjoint zero forcing sets. ∎
@@
If $G=K_{n}$ for $n\geq 2$, then $\operatorname{Z}_{t}(G)=n$.
@@
Let $G=K_{n}$ for some $n\geq 2$. Then $\operatorname{Z}(K_{n})=n-1$ for all $n\geq 2$, and any set of n1n-1 vertices is a minimum zero forcing set. Therefore, the set of all $n$ vertices is a fault tolerant zero forcing set, and thus $\operatorname{Z}_{t}(K_{n})\leq n$. We have $\operatorname{Z}_{t}(K_{n})\geq n$ by 2.5. ∎
@@
$\operatorname{Z}(G)=n-1$, and all minimum zero forcing sets are composed of all but one leaf. Thus, the set of all leaves is a 1-fault tolerant set, so $\operatorname{Z}_{t}(G)\leq n$. We have $\operatorname{Z}_{t}(G)\geq n$ by 2.5. ∎
@@
If $G=W_{n}$ is a wheel on $n\geq 6$ vertices, then $\operatorname{Z}_{t}(G)=5$.
@@
Let $G=W_{n}$ for $n\geq 6$. It was shown in [16] that $\operatorname{Z}(G)=3$. Furthermore, up to isomorphism there are two minimum zero forcing sets, and these are precisely the minimal zero forcing sets. The first minimum zero forcing set is $B_{1}=\{v_{1},v_{2},v_{3}\}$, where $\deg(v_{1})=\deg(v_{2})=3$, $v_{1}v_{2}\in E(G)$, and $\deg(v_{3})=n-1$, while the second is $B_{2}=\{w_{1},w_{2},w_{3}\}$, where $\deg(w_{1})=\deg(w_{2})=\deg(w_{3})=3$ and $w_{1}w_{2},w_{2}w_{3}\in E(G)$. By 2.5, we have $\operatorname{Z}_{t}(G)\geq 4$. We will show there is no $B$ such that $|B|=4$ is a fault tolerant zero forcing set and construct a fault tolerant zero forcing set $B$ such that $|B|=5$.
@@ 4.1 Compatible path covers
Let $G$ be a graph and $\mathcal{P}=\{P_{i}\}_{i=1}^{r}$ a path cover of $G$. Define $L(\mathcal{P})$ to be the endpoints of paths in $\mathcal{P}$, that is $L(\mathcal{P})=\{v\in V(T)\ |\ \deg_{\mathcal{P}}(v)\leq 1\}$. If $\mathfrak{P}$ is a set of path covers, then $L(\mathfrak{P})=\bigcup_{\mathcal{P}\in\mathfrak{P}}L(\mathcal{P})$.
@@
Let $G$ be a graph and $v\in V(G)$. A set $\mathfrak{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},\dots,\mathcal{P}_{r}\}$ of path covers of $G$ is called compatible for $S\subseteq L(\mathfrak{P})$ if for every $v\in S$ both of the following hold:
@@
    for each $u_{i}\in L(\mathcal{P}_{i})$ such that $\deg_{\mathcal{P}_{i}}(v)=1$, there exists $\mathcal{P}_{j}$ such that $\deg_{\mathcal{P}_{j}}(v)=2$ or $\mathrm{comp}_{\mathcal{P}_{j}}(u_{i})\neq\mathrm{comp}_{\mathcal{P}_{j}}(v)$ and $\deg_{\mathcal{P}_{j}}(u_{i})\geq 1$.
While the definition of a compatible collection is quite technical, the idea is that if a collection is compatible for $S$, then the vertices of $S$ can always be forced from two different path covers and is not required to initiate forcing on path. Thus $S$ is not required in a minimal fault tolerant zero forcing set. We begin with a trivial example to show how compatible collections generalize path covers, before exhibiting a more interesting case.
@@
Consider the Graph $E_{1}$ in Figure 4.1. The collection $\mathfrak{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},\mathcal{P}_{3},\mathcal{P}_{4}\}$, where each $\mathcal{P}_{i}$ is given in Figure 4.2, is a set of path covers compatible for $\{1,3,4,5\}$.
@@
Given a compatible collection of path covers $\mathfrak{P}$ for $S$, the irredundant vertices of $\mathfrak{P}$ are $I(\mathfrak{P})=L(\mathfrak{P})\setminus S$.
A collection of path covers $\mathfrak{P}$ of a graph $G$ is a minimum compatible collection for $G$ if $|I(\mathfrak{P})|\leq|I(\mathfrak{Q})|$ for every collection of path covers $\mathfrak{Q}$ of $G$. The cardinality of a minimum compatible collection for $G$ is called the irredundance number of $G$ is denoted by $\mathrm{I}(G)$.
@@
Consider $w\not\in B$. Since $w$ must be forced, we have $\deg_{\mathcal{P}_{v}}(w)\geq 1$ for every $v\in B$. If $w\not\in L(\mathfrak{P})$, then $\deg_{\mathcal{P}_{v}}(w)\geq 2$ for all $v\in B$ and $\mathfrak{P}$ is compatible for $w$. So suppose $w\in L(\mathfrak{P})$. For every $\mathcal{P}_{v}$ such that $\deg_{\mathcal{P}_{v}}(w)=1$, the other endpoint of $\mathrm{comp}_{\mathcal{P}_{v}}(w)$ must be in $B\setminus\{v\}$, and thus $\mathfrak{P}$ is compatible for $w$. Therefore $\mathfrak{P}$ is compatible for every $w\not\in B$, and since $I(\mathfrak{P})=|B|$, we have the result. ∎
@@
In general, determining whether a set of compatible path covers for a graph $G$ is minimum is difficult and is a topic that requires further study. All path covers can be enumerated and thus an algorithm exists to test minimality, but this is not efficient to implement. Some simplifications exist, however, and as an example we offer the next result.
@@
If $\mathfrak{P}$ is a minimum compatible collection for $G$ and $v\in L(\mathfrak{P})$ is isolated in every path cover of $\mathfrak{P}$, then $\mathfrak{P}-v$ is a minimum compatible collection of GvG-v. In particular, under these hypotheses I(G)=I(Gv)+1I(G)=I(G-v)+1